OpenStudy (anonymous):

STAT help please 1. Suppose X has normal distribution with mean 5 and variance 1. Find the probability that 5 will be in the random interval (X-1,X+1). 2. Suppose that X has uniform distribution on the interval [0,1]. Find the probability that 1/2 will be in the random interval (X - 1/4, X + 1/4). 3. Suppose X has normal distribution with mean 5 and variance 1. Find the probability that the random variable 2X will be in the random interval (X-1 ,X+1).

OpenStudy (zarkon):

\[P(X-1<5<X+1)\] \[=P(-1<5-X<1)\] \[=P(-6<-X<-4)\] \[=P(4<X<6)=\cdots\]